# GR-Gaussian: Graph-Based Radiative Gaussian Splatting for Sparse-View CT Reconstruction

### v1

@@ Abstract
Figure 1: Results. We compare our method to two NeRF-based methods NAF(Zha, Zhang, and Li 2022), SAX-NeRF(Cai et al. 2024b)) and $R^{2}$-GS(Zha et al. 2024) in terms of visual quality and PSNR (dB). Our method mitigates needle-like artifacts and achieves superior CT reconstruction quality under sparse-view conditions.
@@ 1 Introduction
Computed Tomography (CT) is a widely used non-invasive imaging technique applied in fields such as medicine, biology, security, and industry. It utilizes X-rays to capture multi-angle 2D projections, which measure the attenuation of rays as they pass through an object(Kak and Slaney 2001). These projections are then processed to reconstruct the 3D density distribution of the object, enabling detailed internal visualization.
In many applications, tomography data acquisition is limited by the number of projections or non-uniform angular sampling. For instance, medical CT reduces the number of projections to minimize X-ray exposure, while electron tomography faces the missing wedge problem due to restricted rotation angles. These constraints make the reconstruction problem highly ill-posed. Traditional analytical methods, such as FBP and FDK(Feldkamp, Davis, and Kress 1984), are computationally efficient but often fail under these conditions, producing artifacts. Model-based optimization approaches(Andersen and Kak 1984; Ma et al. 2024b), which use nonlinear regularizers and iterative solvers (e.g., ART), achieve better results but are computationally expensive.
Deep learning-based methods(Lin et al. 2023; Chung et al. 2023; Liu et al. 2023) demonstrate strong performance but require large labeled datasets and long training times(Ma et al. 2025a, c). They also struggle with out-of-distribution objects. NeRF-based(Mildenhall et al. 2021; Ma et al. 2025b, 2024a) approaches show promise in per-case reconstruction but are extremely time-consuming due to extensive point sampling during volume rendering(Zha, Zhang, and Li 2022; Cai et al. 2024b; Zang et al. 2021). These methods highlight the trade-off between reconstruction quality and computational efficiency.
Recently, 3D Gaussian Splatting (3DGS)(Kerbl et al. 2023; Yu et al. 2024; Guédon and Lepetit 2024) has demonstrated superior quality and efficiency over NeRF in view synthesis tasks(Lu et al. 2024; Liang et al. 2024). However, when applied to sparse-view tomographic reconstruction, 3DGS-based methods exhibit characteristic needle-like artifacts. Through analysis, we identify that these artifacts arise from the retention of large Gaussian kernels with small gradients during the reconstruction process. This limitation stems from the lack of consideration for inter-point relationships, which results in biased density distributions and insufficient splitting of Gaussian kernels.
@@
We validate our framework through extensive experiments on both simulated X-3D dataset(Zha et al. 2024) and Real-world CT datasets (Society 2024), comparing it against multiple baseline methods. The results demonstrate that GR-Gaussian outperforms existing approaches in reconstruction quality, achieving PSNR gains of 0.67 dB and 0.92 dB, and SSIM improvements of 0.011 and 0.021. Our main contributions are summarized as follows:
@@ 2.1 Tomographic Reconstruction
Computed Tomography (CT) is a critical imaging technique widely used in medicine(Hounsfield 1980; Katsuragawa and Doi 2007) and industry(De Chiffre et al. 2014; Zang et al. 2019). Sparse-view CT reconstruction, where the number of projections is limited, poses significant challenges for traditional methods. Traditional analytical methods, such as Filtered Back Projection (FBP) and its 3D counterpart FDK(Feldkamp, Davis, and Kress 1984), are computationally efficient but prone to severe streak artifacts in sparse-view scenarios, leading to degraded image quality. Iterative methods(Andersen and Kak 1984; Sidky and Pan 2008; Manglos et al. 1995; Sauer and Bouman 2002), which optimize reconstruction through regularized energy minimization, can suppress artifacts but are computationally expensive and risk losing fine structural details. Deep learning methods(Lin et al. 2023; Chung et al. 2023; Liu et al. 2023) have shown promise in CT reconstruction, particularly in sparse-view scenarios. Supervised approaches leverage large datasets to learn semantic priors for tasks like projection inpainting(Anirudh et al. 2018; Ghani and Karl 2018) and volume denoising(Chung et al. 2023; Lee et al. 2023; Liu et al. 2023, 2020), but struggle with generalization to unseen data. Self-supervised methods, inspired by NeRF(Mildenhall et al. 2021), optimize density fields using photometric losses(Cai et al. 2024b; Zha, Zhang, and Li 2022; Zang et al. 2021) but are computationally prohibitive due to extensive point sampling during rendering.
@@
3D Gaussian Splatting has demonstrated remarkable performance in RGB tasks, such as surface reconstruction(Li et al. 2025; Lin et al. 2025; Hou et al. 2025), dynamic scene modeling(Zhu et al. 2025; Huang et al. 2025; Wu et al. 2024), human avatar creation(Qu et al. 2025; Li et al. 2024), and 3D generation, outperforming NeRF in terms of speed by leveraging highly parallelized rasterization for image rendering. Recent studies have attempted to extend 3DGS to X-ray imaging. For example, Cai(Cai et al. 2024a) empirically modified 3DGS to synthesize novel-view X-ray projections, primarily using it as a data augmentation tool for traditional algorithms rather than directly generating CT models. Zha(Zha et al. 2024) represented the density field with customized Gaussian kernels, yet they replaced the efficient rasterization with existing CT simulators. While effective in certain cases, their approach overfits sparse-view projections, leading to reconstruction failures.
Figure 2: Training pipeline of GR-Gaussian. (a) Overall training pipeline. (b) Denoised Point Cloud Initialization Strategy. (c) Pixel-Graph-Aware Gradient Strategy.
3 Preliminary
3.1 Radiative Gaussians
To represent the target object, previous approach(Zha et al. 2024) utilize a set of learnable 3D Gaussian kernels, denoted as $\mathbb{G}^{3}=\{G_{i}^{3}\}_{i=1,\cdots,M}$, which we refer to as Radiative Gaussians. Each kernel $G_{i}^{3}$ defines a local Gaussian-shaped density field, expressed as:
@@ 3.1 Radiative Gaussians
3.2 Pixel-Aware Gradient
@@
4 Method
Figure 3: The scanned object is represented as graph-based radiative Gaussians, optimized using real X-ray projections to retrieve the density volume via voxelization.
@@ 4.1 Representing Objects with Graph-based Radiative Gaussians
=$\mathcal{E}=\{(\mathbf{p}_{i},\mathbf{p}_{j})\mid\mathbf{p}_{j}\in\text{KNN}(\mathbf{p}_{i})\text{ and }\mathbf{p}_{i}\in\text{KNN}(\mathbf{p}_{j})\}.$ (9)
@@ 4.2 Training Graph-based Radiative Gaussians
Denoised Point Cloud Initialization Strategy
@@
where $|j|^{2}=j_{1}^{2}+j_{2}^{2}+\cdots+j_{n}^{2}$, and $Z=\sum_{j\in\mathcal{N}(i)}G(j;\sigma_{d})$ is the normalization factor ensuring the kernel weights sum to 1. Empty regions are excluded using a density threshold $\tau$, and $M$ points are randomly sampled as kernel positions. Following (Zha et al. 2024), Gaussian scales are set based on nearest neighbor distances, assuming no rotation. Central densities are directly queried from the denoised volume.
Pixel-Graph-Aware Gradient Strategy
@@
Figure 4: The Pixel-Graph-Aware Gradient Strategy leverages density differences between Gaussian kernels by constructing a graph to encode point-to-point relationships, enhancing gradient computation and enabling effective splitting of large kernels with low gradients.
Density Voxelizer
We employ a voxelizer, denoted as $\mathbf{V}$, to efficiently extract a density volume $V\in R^{X\times Y\times Z}$ from the graph-based radiative representation, defined as $V=\mathbf{V}(\mathcal{G})$. The voxelizer (Zha et al. 2024) divides the target space into $8\times 8\times 8$ 3D tiles and performs Gaussian culling to retain kernels with a 99% confidence of intersecting each tile. Building on this, we incorporate graph-based relationships into the voxelization process. The final voxel values are computed in parallel, as detailed in Eq. 13.
Optimization
To optimize the radiative Gaussians, we employ stochastic gradient descent. The total loss function consists of the photometric loss $\mathcal{L}_{1}$, the D-SSIM loss $\mathcal{L}_{ssim}$(Wang et al. 2004), the 3D total variation (TV) regularization $\mathcal{L}_{tv}$, and the graph Laplacian regularization $\mathcal{L}_{lap}$. The graph Laplacian regularization term $\mathcal{L}_{lap}$ is defined as:
@@
Table 1: Detailed quantitative results under 25-view. Comparison of our GR-Gaussian with different methods on the X-3D dataset and Real-world dataset. HO: Human Organs, BS: Biological Specimens, AO: Artificial Objects (all from the X-3D dataset); RD: Real-world Dataset. We colorize the best, second-best, and third-best numbers.
@@ 5.1 Experimental Settings
Dataset
Following $R^{2}$-Gaussian , we utilize the publicly available X-3D dataset to evaluate our method. This dataset(Zha et al. 2024) encompasses diverse categories, including human organs(chest, foot, head, jaw, and pancreas), artificial objects(backpack, engine, mount, present, and teapot), and biological specimens(beetle, bonsai, broccoli, kingsnake, and pepper), providing a comprehensive benchmark for reconstruction performance. Using the tomography toolbox TIGRE, we generate $512\times 512$ projections along a full circular trajectory, incorporating electric noise and ponton scatter to simulate realistic conditions. Subsequently, CT volumes are scanned within an angular range of $0^{\circ}$ to $360^{\circ}$ to produce 25-view projections. For the Real-World datasets(Society 2024), the projections are directly acquired using CT scanners and correspond to specific angles. Each projection has a resolution of $560\times 560$. To achieve sparse-view conditions, we uniformly select 25 projection angles from the full angular range of $0^{\circ}$ to $360^{\circ}$. This dataset including pine, seashell and walnut.
Implementation Details
The GR-Gaussian framework is implemented using PyTorch(Paszke 2019), PyTorch Geometric(Fey and Lenssen 2019), and CUDA(Guide 2013), and optimized with the Adam optimizer(Diederik 2014). Initial learning rates for location, density, scale, and rotation are set to 0.0002, 0.01, 0.005, and 0.001, respectively, with exponential decay reducing them to 10% of their initial values. Regularization parameters include a television volume level of $D=32$, loss weights $\lambda_{\mathrm{ssim}}=0.25$ and $\lambda_{\mathrm{tv}}=0.05$, a Graph Laplacian weight $\mathcal{L}_{\mathrm{lap}}=8\times 10^{-4}$, a Gaussian denoising parameter $\sigma_{d}=3$, and a gradient computation weight $\lambda_{\mathrm{g}}=1\times 10^{-4}$. The framework uses $M=50,000$ Gaussians, a density threshold $\tau=0.001$, and $k=6$ nearest neighbors. A dynamic stopping criterion ($Iter_{stop}$) evaluates PSNR every 500 iterations, terminating if PSNR decreases by more than 0.5%. All experiments were conducted on an RTX 4090 GPU. Quantitative evaluation employed PSNR and SSIM(Wang et al. 2004), with PSNR assessing 3D reconstruction accuracy and SSIM evaluating structural consistency across 2D slices.
Figure 5: The left shows CT reconstructions from the X-3D dataset, covering three categories: chest, bonsai, and teapot. The right displays Real-world dataset reconstructions with colorized slices to highlight details, all under 25-view conditions.
@@ 5.2 Results and Evaluation
The experiments aim to evaluate the performance of GR-Gaussian in sparse-view tomographic reconstruction, focusing on reconstruction quality across X-3D and Real-world datasets. On the X-3D dataset, GR-Gaussian outperforms the SOTA 3DGS-based method (Zha et al. 2024), NeRF-based methods (Zha, Zhang, and Li 2022; Cai et al. 2024b) and three traditional methods (Feldkamp, Davis, and Kress 1984; Andersen and Kak 1984; Sidky and Pan 2008) in both PSNR and SSIM. As show in Table 1, it achieves a PSNR increase of 0.67 dB and an SSIM improvement of 0.011, demonstrating its ability to suppress artifacts and enhance structural consistency. In the experiments with real-world datasets, GR-Gaussian demonstrates significant robustness against noise and artifacts, achieving a PSNR improvement of 0.92 dB and an SSIM improvement of 0.021. These results highlight the method’s adaptability to practical scenarios, where data imperfections pose significant challenges to reconstruction quality. As shown in Figure 5, GR-Gaussian produces visually superior reconstructions with sharper edges and fewer streak artifacts compared to other methods. These qualitative improvements align with the quantitative results, further validating the effectiveness of the proposed framework.
@@
Figure 6: Ablation study results highlight the impact of PGA and De-Init in enhancing reconstruction quality.
@@ 5.3 Ablation Study
k 4 5 6 7 8
X-3D PSNR 35.80 35.84 35.86 35.87 35.89
@@
Real World PSNR 35.82 35.87 35.95 35.97 36.01
@@
Real World PSNR 35.90 35.93 35.95 35.91 35.85
@@
Table 3: Performance comparison for different k-nearest neighbors and $\sigma_{d}$. Analysis of the X-3D dataset and Real-world dataset.
Parameter Analysis
@@
Figure 7: Iteration Analysis Reconstruction results of GR-Gaussian across different iterations, illustrating the impact of iteration count on PSNR.
Iteration Analysis

### v2

@@ Abstract
Fig 1: Visual results of GR-Gaussian. We compare GR-Gaussian to two NeRF-based methods NAF[44], SAX-NeRF[4]) and $R^{2}$-GS[45] in terms of visual quality and PSNR (dB). Our method mitigates needle-like artifacts and achieves superior CT reconstruction quality under sparse-view conditions.
footnotetext: * Equal contribution.footnotetext: $\dagger$ Corresponding Authors.
@@ 1 Introduction
Fig 2: Training pipeline of GR-Gaussian. (a) Overall training pipeline. (b) Denoised Point Cloud Initialization Strategy. (c) Pixel-Graph-Aware Gradient Strategy.
ake the reconstruction problem highly ill-posed. Traditional analytical methods, such as FBP and FDK[8], are computationally efficient but often fail under these conditions, producing artifacts. Model-based optimization approaches[1, 28], which use nonlinear regularizers and iterative solvers (e.g., ART), achieve better results but are computationally expensive.
Deep learning-based methods[23, 5, 24] demonstrate strong performance but require large labeled datasets and long training times[29, 31]. They also struggle with out-of-distribution objects. NeRF-based[33, 30, 27] approaches show promise in per-case reconstruction but are extremely time-consuming due to extensive point sampling during volume rendering[44, 4, 43]. These methods highlight the trade-off between reconstruction quality and computational efficiency.
Recently, 3D Gaussian Splatting (3DGS)[17, 41, 11] has demonstrated superior quality and efficiency over NeRF in view synthesis tasks[26, 21]. However, when applied to sparse-view tomographic reconstruction, 3DGS-based methods exhibit characteristic needle-like artifacts. Through analysis, we identify that these artifacts arise from the retention of large Gaussian kernels with small gradients during the reconstruction process. This limitation stems from the lack of consideration for inter-point relationships, which results in biased density distributions and insufficient splitting of Gaussian kernels.
@@
We validate our framework through extensive experiments on both simulated X-3D dataset[45] and Real-world CT datasets [38], comparing it against multiple baseline methods. The results demonstrate that GR-Gaussian outperforms existing approaches in reconstruction quality, achieving PSNR gains of 0.67 dB and 0.92 dB, and SSIM improvements of 0.011 and 0.021. Our main contributions are summarized as follows:
@@
Fig 3: The scanned object is represented as graph-based radiative Gaussians, optimized using real X-ray projections to retrieve the density volume via voxelization.
@@ 2.1 Tomographic Reconstruction
Computed Tomography (CT) is a critical imaging technique widely used in medicine[14, 16] and industry[6, 42]. Sparse-view CT reconstruction, where the number of projections is limited, poses significant challenges for traditional methods. Traditional analytical methods, such as Filtered Back Projection (FBP) and its 3D counterpart FDK[8], are computationally efficient but prone to severe streak artifacts in sparse-view scenarios, leading to degraded image quality. Iterative methods[1, 37, 32, 36], which optimize reconstruction through regularized energy minimization, can suppress artifacts but are computationally expensive and risk losing fine structural details. Deep learning methods[23, 5, 24] have shown promise in CT reconstruction, particularly in sparse-view scenarios. Supervised approaches leverage large datasets to learn semantic priors for tasks like projection inpainting[2, 10] and volume denoising[5, 18, 24, 25], but struggle with generalization to unseen data. Self-supervised methods, inspired by NeRF[33], optimize density fields using photometric losses[4, 44, 43] but are computationally prohibitive due to extensive point sampling during rendering.
@@
3D Gaussian Splatting has demonstrated remarkable performance in RGB tasks, such as surface reconstruction[19, 22, 13], dynamic scene modeling[46, 15, 40], human avatar creation[35, 20], and 3D generation, outperforming NeRF in terms of speed by leveraging highly parallelized rasterization for image rendering. Recent studies have attempted to extend 3DGS to X-ray imaging. For example, Cai[3] empirically modified 3DGS to synthesize novel-view X-ray projections, primarily using it as a data augmentation tool for traditional algorithms rather than directly generating CT models. Zha[45] represented the density field with customized Gaussian kernels, yet they replaced the efficient rasterization with existing CT simulators. While effective in certain cases, their approach overfits sparse-view projections, leading to reconstruction failures.
3 Method
3.1 Preliminaries
3.1.1 Radiative Gaussians
To represent the target object, previous approach[45] utilize a set of learnable 3D Gaussian kernels, denoted as $\mathbb{G}^{3}=\{G_{i}^{3}\}_{i=1,\cdots,M}$, which we refer to as Radiative Gaussians. Each kernel $G_{i}^{3}$ defines a local Gaussian-shaped density field, expressed as:
@@ 3.1.1 Radiative Gaussians
3.1.2 Pixel-Aware Gradient
@@ 3.2 Representing Objects with Graph-based Radiative Gaussians
$\mathcal{E}=\{(\mathbf{p}_{i},\mathbf{p}_{j})\mid\mathbf{p}_{j}\in\text{KNN}(\mathbf{p}_{i})\text{ and }\mathbf{p}_{i}\in\text{KNN}(\mathbf{p}_{j})\}.$ (9)
@@ 3.3 Training Graph-based Radiative Gaussians
3.3.1 Denoised Point Cloud Initialization Strategy
@@
where $|j|^{2}=j_{1}^{2}+j_{2}^{2}+\cdots+j_{n}^{2}$, and $Z=\sum_{j\in\mathcal{N}(i)}G(j;\sigma_{d})$ is the normalization factor ensuring the kernel weights sum to 1. Empty regions are excluded using a density threshold $\tau$, and $M$ points are randomly sampled as kernel positions. Following [45], Gaussian scales are set based on nearest neighbor distances, assuming no rotation. Central densities are directly queried from the denoised volume.
3.3.2 Pixel-Graph-Aware Gradient Strategy
@@
Fig 4: The Pixel-Graph-Aware Gradient Strategy leverages density differences between Gaussian kernels by constructing a graph to encode point-to-point relationships, enhancing gradient computation and enabling effective splitting of large kernels with low gradients.
3.3.3 Density Voxelizer
We employ a voxelizer, denoted as $\mathbf{V}$, to efficiently extract a density volume $V\in R^{X\times Y\times Z}$ from the graph-based radiative representation, defined as $V=\mathbf{V}(\mathcal{G})$. The voxelizer [45] divides the target space into $8\times 8\times 8$ 3D tiles and performs Gaussian culling to retain kernels with a 99% confidence of intersecting each tile. Building on this, we incorporate graph-based relationships into the voxelization process. The final voxel values are computed in parallel, as detailed in Eq. 13.
3.3.4 Optimization
To optimize the radiative Gaussians, we employ stochastic gradient descent. The total loss function consists of the photometric loss $\mathcal{L}_{1}$, the D-SSIM loss $\mathcal{L}_{ssim}$[39], the 3D total variation (TV) regularization $\mathcal{L}_{tv}$, and the graph Laplacian regularization $\mathcal{L}_{lap}$. The graph Laplacian regularization term $\mathcal{L}_{lap}$ is defined as:
@@ 3.3.4 Optimization
Table 1: Detailed quantitative results under 25-view. Comparison of our GR-Gaussian with different methods on the X-3D dataset and Real-world dataset. HO: Human Organs, BS: Biological Specimens, AO: Artificial Objects (all from the X-3D dataset); RD: Real-world Dataset.
@@ 4.1 Experimental Settings
4.1.1 Dataset
Following $R^{2}$-Gaussian , we utilize the publicly available X-3D dataset to evaluate our method. This dataset[45] encompasses diverse categories, including human organs(chest, foot, head, jaw, and pancreas), artificial objects(backpack, engine, mount, present, and teapot), and biological specimens(beetle, bonsai, broccoli, kingsnake, and pepper), providing a comprehensive benchmark for reconstruction performance. Using the tomography toolbox TIGRE, we generate $512\times 512$ projections along a full circular trajectory, incorporating electric noise and ponton scatter to simulate realistic conditions. Subsequently, CT volumes are scanned within an angular range of $0^{\circ}$ to $360^{\circ}$ to produce 25-view projections. For the Real-World datasets[38], the projections are directly acquired using CT scanners and correspond to specific angles. Each projection has a resolution of $560\times 560$. To achieve sparse-view conditions, we uniformly select 25 projection angles from the full angular range of $0^{\circ}$ to $360^{\circ}$. This dataset including pine, seashell and walnut.
4.1.2 Implementation Details
The GR-Gaussian framework is implemented using PyTorch[34], PyTorch Geometric[9], and CUDA[12], and optimized with the Adam optimizer[7]. Initial learning rates for location, density, scale, and rotation are set to 0.0002, 0.01, 0.005, and 0.001, respectively, with exponential decay reducing them to 10% of their initial values. Regularization parameters include a television volume level of $D=32$, loss weights $\lambda_{\mathrm{ssim}}=0.25$ and $\lambda_{\mathrm{tv}}=0.05$, a Graph Laplacian weight $\mathcal{L}_{\mathrm{lap}}=8\times 10^{-4}$, a Gaussian denoising parameter $\sigma_{d}=3$, and a gradient computation weight $\lambda_{\mathrm{g}}=1\times 10^{-4}$. The framework uses $M=50,000$ Gaussians, a density threshold $\tau=0.001$, and $k=6$ nearest neighbors. A dynamic stopping criterion ($Iter_{stop}$) evaluates PSNR every 500 iterations, terminating if PSNR decreases by more than 0.5%. All experiments were conducted on an RTX 4090 GPU. Quantitative evaluation employed PSNR and SSIM[39], with PSNR assessing 3D reconstruction accuracy and SSIM evaluating structural consistency across 2D slices.
Fig 5: The left shows CT reconstructions from the X-3D dataset, covering three categories: chest, bonsai, and teapot. The right displays Real-world dataset reconstructions with colorized slices to highlight details, all under 25-view conditions.
@@ 4.2 Results and Evaluation
The experiments aim to evaluate the performance of GR-Gaussian in sparse-view tomographic reconstruction, focusing on reconstruction quality across X-3D and Real-world datasets. On the X-3D dataset, GR-Gaussian outperforms the SOTA 3DGS-based method [45], NeRF-based methods [44, 4] and three traditional methods [8, 1, 37] in both PSNR and SSIM. As show in Table 1, it achieves a PSNR increase of 0.67 dB and an SSIM improvement of 0.011, demonstrating its ability to suppress artifacts and enhance structural consistency. In the experiments with real-world datasets, GR-Gaussian demonstrates significant robustness against noise and artifacts, achieving a PSNR improvement of 0.92 dB and an SSIM improvement of 0.021. These results highlight the method’s adaptability to practical scenarios, where data imperfections pose significant challenges to reconstruction quality. As shown in Figure 5, GR-Gaussian produces visually superior reconstructions with sharper edges and fewer streak artifacts compared to other methods. These qualitative improvements align with the quantitative results, further validating the effectiveness of the proposed framework.
@@
Fig 6: Ablation study results highlight the impact of PGA and De-Init in enhancing reconstruction quality.
@@ 4.3.1 Component Analysis
Table 3: Quantitative results for different k-nearest neighbors and $\sigma_{d}$. Analysis of the X-3D dataset and Real-world dataset.
Metric k=4 k=5 k=6 k=7 k=8
X-3D Dataset
PSNR 35.80 35.84 35.86 35.87 35.89
@@ X-3D Dataset
Real World Dataset
PSNR 35.82 35.87 35.95 35.97 36.01
@@
Real World Dataset
PSNR 35.90 35.93 35.95 35.91 35.85
@@
4.3.2 Parameter Analysis
@@
Table 4: Quantitative results of SSGU extension.
Method w/o SSGU w/ SSGU $\lambda=3$ w/ SSGU $\lambda=5$
SDS 40.37s 14.12s 8.62s
DDS 66.89s 22.65s 13.90s
CoSD 344.76s 128.97s 79.15s
Fig 7: Iteration Analysis Reconstruction results of GR-Gaussian across different iterations, illustrating the impact of iteration count on PSNR.
4.3.3 Iteration Analysis